OpenStudy (ttp):

Solve 2x2 + 3x + 6 = 0. Round solutions to the nearest hundredth. x ≈ -2.64 and x ≈ 1.14 x ≈ -1.14 and x ≈ 2.64 No real solutions x ≈ -1.11 and x ≈ -4.89

OpenStudy (anonymous):

no real solutions

OpenStudy (anonymous):

[-b-+sqrt(b^2-4ac)]/(2a)

OpenStudy (anonymous):

sqrt(3^2-4*2*6)=sqrt(9-48)=sqrt(-39)

OpenStudy (anonymous):

no real solution